OpenStudy (anonymous):

The magnitude of an earthquake is found by the equation M=2/3 log E/10^118, where M is the magnitude and E is the energy released. Find the magnitude of an earthquake that released 10^203 ergs of energy.

OpenStudy (valpey):

\[M = \frac{2}{3} \log{\frac{10^{203}}{10^{118}}} \] Is this correct?

OpenStudy (valpey):

And that is a base 10 log, right, not a natural log?

OpenStudy (anonymous):

yeah, I just need help solving it

OpenStudy (valpey):

Okay, can you simplify\(\frac{10^{203}}{10^{118}}\)?

OpenStudy (anonymous):

no?...

OpenStudy (valpey):

How about \(\frac{10^3}{10^1}\), what does that equal?

OpenStudy (anonymous):

100

OpenStudy (valpey):

Right, 100 is \(10^2\) or \(10^{3-1}\) so subtract the powers.

OpenStudy (anonymous):

ok so it'd be 10^85

OpenStudy (valpey):

Cool. So what is \(\log_{10}{10^{85}}\)?

OpenStudy (anonymous):

85

OpenStudy (valpey):

You good from here?

OpenStudy (anonymous):

yes tysm

OpenStudy (valpey):

Cheers